OpenStudy (anonymous):

y=4/5x-38/5 Find x when y=7

jimthompson5910 (jim_thompson5910):

I'll start you off y=4/5x-38/5 7=4/5x-38/5 5*7 = 5*[ 4/5x-38/5 ] 35 = 4x - 38 and let you finish up

OpenStudy (anonymous):

Okay thats what I needed. I didn't know what to do with the fractions ha thank you!

jimthompson5910 (jim_thompson5910):

yeah you could deal with the fractions throughout the entire solving process or just multiply both sides by the LCD (in this case 5) to clear out the fractions

OpenStudy (anonymous):

I got 73/4 thanks!

jimthompson5910 (jim_thompson5910):

x = 73/4 is correct